OpenStudy (anonymous):

5. Eliminate the parameter and write the corresponding rectangular equations: x=3-2sin theta and y=1+5cos theta

OpenStudy (anonymous):

(x-3)/2=-sin theta (y-1)/5=cos theta (x-3)^2/4 = sin^2 theta (y-1)^2/25 = cos^2 theta adding them bouth: (x-3)^2/4 + (y-1)^2/25=1

OpenStudy (anonymous):

Thank you so much!! :D

OpenStudy (anonymous):

Another quick question: Graph the equation r= 1 – cosФ. Use observation to state the symmetry

OpenStudy (anonymous):

 it's something like that